OpenStudy (anonymous):

2x^4-4x^2-48. (a) Find all the real zeros of the polynomial function. b) Determine the multiplicity of each zero and the number of turning points of the graph of the function. (c) The number of turning points.

OpenStudy (amistre64):

step one is to factor out the commons

OpenStudy (amistre64):

its also best to see this as a quadratic in disguise; say u=x^2 and rewrite

OpenStudy (amistre64):

2u^2-4u -48 2 (u^2-2u -12)

OpenStudy (amistre64):

now its pretty much a normal factoring or the quad formula

OpenStudy (amistre64):

1/2 * 2 +- sqrt(100) so its u = 6 or u = 4 i think

OpenStudy (amistre64):

-12 up there should be -24

OpenStudy (amistre64):

u = x^2 ; and u = 6 or 4 6 = x^2 x = +- sqrt(6) 4 = x^2; x = +- 2

OpenStudy (amistre64):

tag and test

OpenStudy (anonymous):

So do I plug in -/+ 2 and -/+ sqrt(6) into the equation?

OpenStudy (amistre64):

to test? yes. since that is what we found it would be good to test them to see if they actually work